OpenStudy (anonymous):

Find the length and width of the rectangle if the area is 9 m square, the width is (x-3) and the length is (x+5). Recall the formula for area of a rectangle is Area = LW.

OpenStudy (anonymous):

9 = (x - 3)(x + 5) 9 = x^2 + 2x - 15 0 = x^2 + 2x - 24 0 = (x + 6)(x - 4) x = -6, 4 Plug it back into the equation. 4 - 3 = 1 4 + 5 = 9 The length is 9 and the width is 1.

OpenStudy (anonymous):